Given- p=1000-50Q-2(Q*2)

a)find TR max level Q

b)Find TR nax level p

C)find TR max

D)find MR when Q=2

e)find AR when Q=3


1
Expert's answer
2021-04-28T09:30:17-0400

a) TR=pQ=(1000-50Q-2Q^2)× Q.

To find TR max we should solve equation TR'=0. So, 1000-2×50Q-2×3Q^2=0; 1000-100Q-6Q^2=0; Q=7

b) p=1000-50×7-2×7^2=1000-350-98=552

c) TRmax=552×7=3864

d)MR=TR'=1000-2×50Q-2×3×Q^2=1000-100Q-6Q^2. When Q=2, MR=1000-200-24=776.

e)AR equals to TR÷Q=p. So, when Q=3, AR=1000-150-18=832.